OpenStudy (anonymous):

Find the sum of the angle measures in an 11-gon?

OpenStudy (saifoo.khan):

use the formula.

OpenStudy (anonymous):

umm whats the formula ;0

OpenStudy (saifoo.khan):

\[{(x-2) \times 180}\]

OpenStudy (anonymous):

sum =180*(n-2) , n = number of sides

OpenStudy (anonymous):

thankzz ;))

OpenStudy (saifoo.khan):

welkzz ;))